OpenStudy (anonymous):

(cosx-1/sinx)- sinx/ cosx-1

OpenStudy (anonymous):

\[\frac{ \cos x-1 }{ sinx }-\frac{ sinx }{ cosx -1 }\]

OpenStudy (abb0t):

combine everything into one fraction.

OpenStudy (anonymous):

make them have the same denominator

OpenStudy (anonymous):

\[\frac{ (cosx -1)^2 -\sin^2x }{ sinx (cosx-1) }\]

OpenStudy (anonymous):

(cosx-1)(cosx-1)(/sinx(cosx-1)- sinx(sinx)/ cosx-1(sinx)

OpenStudy (anonymous):

combine what?

OpenStudy (anonymous):

got it? or need more explanation? I ll come up on another way.

OpenStudy (anonymous):

now, (a - b ) (a - b) = (a - b)^2 yours (cosx -1)(cosx -1) = (cosx -1 )^2 got it?

OpenStudy (anonymous):

so, I just combine to have only 1 fraction. I can do it because I have the same denominator already (the first step I did) got it?

OpenStudy (anonymous):

if you cannot, don't try. just follow me and say yes/no if you get or not.

OpenStudy (anonymous):

so, are you with me now?

OpenStudy (anonymous):

continue. you have (cos x-1)^2 - sin^2x at numerator, right? open the square. do it at the paper, I do it here, you just compare ours

OpenStudy (anonymous):

\[\frac{ \cos^2x -2cosx + 1-\sin^2x}{sinx (cosx -1) }\]

OpenStudy (anonymous):

replace 1 = sin^2x + cos^2 x

OpenStudy (anonymous):

you have

OpenStudy (anonymous):

\[\frac{ \cos^2x-2cosx+\sin^2x+\cos^2x-\sin^2x }{ sinx(cosx-1)}\]

OpenStudy (anonymous):

from the sin^2x +cos^2x=1 but this one is negative

OpenStudy (anonymous):

cancel out the sin^2

OpenStudy (anonymous):

you mean replace it with 1-cos^x?

OpenStudy (anonymous):

replace 1 only, see mine, again

OpenStudy (anonymous):

what did you replaced it with? I though 1-cos^2x

OpenStudy (anonymous):

It is (cos x -1)^2 not yours

OpenStudy (anonymous):

yes but I didn't see it in your answer

OpenStudy (anonymous):

no I wrote it according to the formula that says: 1-cos^2x= sin^2x

OpenStudy (anonymous):

i didn't give out the answer, we are on the way

OpenStudy (anonymous):

i meant i replaced it with that

OpenStudy (anonymous):

I make the stuff base on what you post. I don't care and I don't know where do you you pick it up.

OpenStudy (anonymous):

i know but i don't get what you had on top since I replaced sin^2x by 1-cos^2x

OpenStudy (anonymous):

i don't replace sin^2 x . It's there still. I just combine terms together

OpenStudy (anonymous):

\[\frac{ 2\cos^2x -2cosx }{ sinx(cosx-1) }= \frac{ 2cosx(cosx-1) }{sinx( cosx -1) }\] = \[\huge\frac{ 2cosx }{ sinx}= 2 cot x\]

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

give me a sec.

OpenStudy (anonymous):

no I am not but let me go over it again

OpenStudy (anonymous):

Madina \[\sin^2x + \cos^2 x =1\]

OpenStudy (anonymous):

This is GREAT i guess i can do whatever works. My moral is up again. Thanks.

OpenStudy (anonymous):

ok, I have to say, thanks god, too

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

the part I cancel out the (cosx-1)?

OpenStudy (anonymous):

just cancel, numerator and denominator, cancel out , done